# Tensor Attention Training: Provably Efficient Learning of Higher-order Transformers

### v2

@@ 4 Exact Tensor Attention Gradient Computation and Complexity
Define the function $\mathsf{F}(x)\in\mathbb{R}^{n\times n^{2}}$ as in Definition D.6. Suppose that $A_{1},A_{2},A_{3}\in\mathbb{R}^{n\times d}$ are three given matrices. Suppose that $\mathsf{Loss}(x)$ is defined as Definition 3.8, where $x=\operatorname{vec}(X)$. Then, we have
@@
Note that, $\mathsf{F}(x)$ is a size $n\times n^{2}$ matrix which is the bottleneck obstacle in time complexity.
@@
Note that ${\cal T}_{\mathrm{mat}}(n,d^{2},n^{2})\geq\Omega(n^{3})$. Thus, the naive tensor attention gradient computation is a complexity obstacle in practice as discussed in Section 1. Based on the closed formulation in Lemma 4.1, we derive our acceleration method, which will be introduced in the following section.
@@ 5 Fast Tensor Attention Gradient Computation
1:procedure FastTensorAttention($A_{1},A_{2},A_{3},A_{4},A_{5},E\in\mathbb{R}^{n\times d},X_{1},X_{2},X_{3},Y_{1},Y_{2}\in\mathbb{R}^{d\times d},n\in\mathbb{N}_{+},d\in\mathbb{N}_{+},\epsilon\in(0,0.1)$) $\triangleright$ Definition 3.9, Theorem 5.2
@@
13:     $U_{4}\leftarrow\operatorname{diag}(\widetilde{\mathsf{R}}(x))U_{1}$triangleright $O(n^{1+o(1)})$ time
14:     $V_{4},~{}W_{4}\leftarrow V_{1},~{}W_{1}$triangleright $O(n^{1+o(1)})$ time
@@
16:     $U_{5},~{}V_{5},~{}W_{5}\leftarrow[U_{3},-U_{4}],~{}[V_{3},V_{4}],~{}[W_{3},W_{4}]$triangleright $O(n^{1+o(1)})$ time
@@
19:     $\widetilde{g}\leftarrow(A_{1}^{\top}U_{5})\odot(A_{2}^{\top}V_{5})\odot(A_{3}^{\top}W_{5})$triangleright $\odot$ in Definition C.3. $O(d^{3}n^{o(1)})$ time
20:     return $\widetilde{g}$triangleright As $d=O(\log n)$, the total complexity is $O(n^{1+o(1)})$ time
@@ 5.1 Main Results for Fast Gradient Computation
Theorem 5.2 (Main result for fast gradient computation).
Assuming the entries of $A_{1},A_{2},$ $A_{3},$ $A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ and $X_{1},X_{2},X_{3},Y_{1},Y_{2}\in\mathbb{R}^{d\times d}$ are represented using $O(\log n)$ bits. Then, there exist an algorithm (Algorithm 1) that runs in $n^{1+o(1)}$ time to solve $\mathsf{ATAttLGC}(n,d=O(\log n),B=o(\sqrt[3]{\log n}),\epsilon=1/\operatorname{poly}(n))$ (see Definition 3.9), i.e., our algorithm computes a gradient matrix $\widetilde{g}\in\mathbb{R}^{d\times d^{2}}$ satisfying $\|\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}-\widetilde{g}\|_{\infty}\leq 1/\operatorname{poly}(n)$.
@@
The assumption in Theorem 5.2 is practical. In practice, especially in recent long context tasks, the $n$ is large, e.g, $n=2\times 10^{6}$ for Google’s Gemini 1.5 Pro  [35], while the model training uses a half-precision floating-point format, e.g., the bit number is $16$. Furthermore, our assumption is “tight”, where if we slightly weaken the assumption, there is no algorithm that can solve the tensor attention gradient computation in truly sub-cubic complexity (Theorem 6.3).
@@ 5.2 Tensor Operation Analysis Techniques
We first introduce a swap rule and a distributed rule, where both proofs are in Appendix C.2.
Fact 5.4 (Swap rule for tensor and matrix product).
Let $W_{1},W_{2}\in\mathbb{R}^{d\times d},A_{1},A_{2}\in\mathbb{R}^{n\times d}$. We have
Fact 5.4 tells us that we can swap the order of tensor operation and matrix multiplication, allowing us to always compute the low dimension first to reduce the complexity.
Fact 5.5.
@@
Fact 5.5 tells us that the multiple tensor operation can be distributed to a different format. If we have some low-rank matrix/tensor, we can distribute them into each component, so that each component can be accelerated via the low-rank property. Intuitively, this allows us to borrow some low-rank benefits from other terms to fix the bottleneck terms.
We provide an important tool whose proof is in Appendix C.2.
@@
Fact 5.7 (Tensor-trick).
@@
Technical novelty over previous works. We generalize beyond the results of AS24b [15], which only provide methods for tensor attention forward. Our paper presents a detailed analysis for tensor attention backward, providing both upper bound and lower bound. Though we build on some results from AS24b [15] and AS24a [14], generalizing to tensor attention backward posed many technical challenges. These challenges are unique to our setting and not presented in previous settings like matrix attention [13, 14] or tensor attention forward [15]. To be more specific, we prove many key properties for tensor operation needed for backward though not needed for forward, including Facts 5.4 (swap rule for tensor and matrix product), 5.5 (distribution rule for tensor and matrix product), C.11 (tensor computation reduction to matrix product), C.12 (distribution rule for tensor computation), Claim C.20 (tensor product to matrix product). Fact 5.4, used as a key part to prove Lemmas F.1 and F.3, gives the swap rule for tensor operations. Lemma 5.6 supports the proof of Fact 5.5 and helps bypass the $O(n^{3}d^{2})$ time complexity bottleneck in the fast computation of $U_{2}$. Fact 5.5, crucial in proving Lemma F.5, shows the distributive nature of tensor operations. Using Facts C.11, C.12, and Claim C.20, we leverage the structure of low-rank matrices $U_{5},V_{5},W_{5}$ to prove Theorem 5.2.
@@ 6.1 $\mathsf{SETH}$ and Tensor Attention Forward Hardness
Hypothesis 6.1 (Strong Exponential Time Hypothesis ($\mathsf{SETH}$), [46]).
@@
Assuming $\mathsf{SETH}$, for any constant $\delta>0$, no algorithm can solve $\mathsf{ATAttC}(n,d=\Theta(\log n),B=\Theta(\sqrt[3]{(1+\gamma)\log n}),\epsilon=n^{\gamma-O(1)})$ (Definition 3.7) in $O(n^{3-\delta})$ time, even if the inputs meet the following conditions for any $\gamma\geq 0$: (1) $V\in\{0,1\}^{n^{2}\times d}$, (2) There exists $B_{a}\leq O((1+\gamma)\log^{2}n)=O(d(\sqrt[3]{(1+\gamma)\log n})^{3})$ where all entries of $Q(K_{1}\oslash K_{2})^{\top}$ are within the range $[1,B_{a}]$ and more than half entries in each row of $Q(K_{1}\oslash K_{2})^{\top}$ are equal to $B_{a}$.
This result shows that assuming $\mathsf{SETH}$, if we just slightly weaken the assumption from $B=O(\sqrt[3]{\log n})$ to $B=\Theta(\sqrt[3]{(1+\gamma)\log n})$ with $\gamma=\omega(1)$, then the tenor attention forward computation is hard, i.e., no algorithm can solve it in truly sub-cubic time.
@@ 6.2 Main Result for Hardness
Theorem 6.3 (Main result for hardness).
@@ 7 Discussion and Conclusion
In this work, we proved that the backward gradient of tensor attention training can be computed in almost linear $n^{1+o(1)}$ time, the same complexity as its forward computation, under a bounded entries assumption. We provided a closed-form solution for the gradient and proposed a fast computation method utilizing polynomial approximation and tensor algebraic techniques. Furthermore, we proved the necessity and tightness of our assumption through hardness analysis, showing that slightly weakening it renders the tensor attention gradient problem unsolvable in truly subcubic time. Our theoretical results establish the feasibility of efficient higher-order transformer training and may facilitate practical applications of tensor attention architectures. Due to space limits, we provide our further discussion and extension in Appendix A. Future work can perform empirical evaluations of the method in practical large language models, and explore how these findings can be implemented in real-world scenarios to enable the development of powerful higher-order models.
Acknowledgement
Research is partially supported by the National Science Foundation (NSF) Grants 2023239-DMS, CCF-2046710, and Air Force Grant FA9550-18-1-0166.
@@ Acknowledgement
In Section A, we provide a further discussion and extension of this work. In Section B, we provide more related works. In Section C, we provide general definitions and several basic facts. In Section D, we show how we calculate the gradient of the loss function. In Section E, we show the time complexity of our algorithm. In Section F, we show that our algorithm can be computed in polynomial time. In Section G, we show the hardness of our algorithm.
@@ Appendix A Further Discussion and Extension
Societal impacts.
We delve into and offer a deeper understanding of the attention mechanism, introducing a novel approach to integrate multi-modality into attention through the tensor attention algorithm. We also demonstrate that the computation of both forward and backward tensor attention can be achieved with almost linear time complexity. Regarding the negative societal impact, since our work is completely theoretical in nature, we do not foresee any potential negative societal impacts which worth pointing out.
Appendix B More Related Work
@@ Appendix C Tensor Operation Background
In Section C.1, we define the notation of computational time and the tensor operation. In Section C.2, we provide some helpful facts of tensor operation. In Section C.3, we provide some helpful facts of vectorization operation. In Section C.4, we provide some helpful facts about the tensor product.
@@ C.1 General definitions and tensor operation
Definition C.2 (Third mode tensor product (,,)($\cdot$,$\cdot$,$\cdot$)( ⋅ , ⋅ , ⋅ )).
@@ C.2 Facts for tensor operation
In Fact C.5, due to definition $V_{1}$ and $V_{2}$ need to have the same number of rows. $W_{1}$ and $W_{2}$ also need to have the same number of rows. $V_{1}$ and $W_{1}$ need to have same number of columns, and $V_{2}$ and $V_{2}$ need to have same number of columns.
Fact C.7 (Swap rule for tensor product and matrix product, Restatement of Fact 5.4).
@@
Fact C.8 (Restatement of Fact 5.5).
@@
where the first step follows from Definition C.2, the second step follows from the property of identity tensor $(\mathsf{I}_{d})_{i,j,k}$, which equals $1$ only when $i=j=k$ and 00 elsewhere, and the last step follows from Definition C.3. ∎
@@
where the first step follows from definition of $C\in\mathbb{R}^{d\times d}$, the second step follows from the matrix can written as the summation of $n_{1}n_{2}$ rank-$1$ matrices, the third step follows from changing the index, the forth step follows from $\underbrace{A_{i+(j-1)n_{1},*}}_{d_{1}\times 1}=\underbrace{a_{1,i}}_{d_{1}\times 1}\circ\underbrace{a_{2,j}}_{d_{1}\times 1}$ by Definition 3.2.
@@ C.3 Facts for vectorization operation
$\operatorname{vec}(ab^{\top})=$ $~{}\operatorname{vec}(\begin{bmatrix}a_{1}b^{\top}\\ a_{2}b^{\top}\\ \dots\\ a_{n}b^{\top}\end{bmatrix})$
@@
Fact C.16 (Tensor-trick, Restatement of Fact 5.7).
@@ C.4 Facts for tensor product
Let $\mathsf{I}_{d}\in\mathbb{R}^{d\times d\times d}$ and $A_{1},A_{2},A_{3}\in\mathbb{R}^{n\times d}$. We have $\mathrm{mat}(\mathsf{I}_{d}(A_{1},A_{2},A_{3}))=A_{1}\mathrm{mat}(\mathsf{I}_{d})(A_{2}\otimes A_{3})^{\top}=A_{1}(A_{2}\oslash A_{3})^{\top}\in\mathbb{R}^{n\times n^{2}}$

### v3

@@ 3 Exact Tensor Attention Gradient Computation and Complexity
Define the function $\mathsf{F}(x)\in\mathbb{R}^{n\times n^{2}}$ as in Definition D.6 (see Fig. 4 for an illustration). Suppose that $A_{1},A_{2},A_{3}\in\mathbb{R}^{n\times d}$ are three given matrices. Suppose that $\mathsf{Loss}(x)$ is defined as Definition 2.8, where $x=\operatorname{vec}(X)$. Then, we have
@@
Note that $\mathsf{F}(x)$ is a size $n\times n^{2}$ matrix which is the bottleneck obstacle in time complexity.
@@
Note that ${\cal T}_{\mathrm{mat}}(n,d^{2},n^{2})\geq\Omega(n^{3})$. Thus, the naive tensor attention gradient computation is a complexity obstacle in practice, as discussed in Section 1. Based on the closed formulation in Lemma 3.1, we derive our acceleration method, which will be introduced in the following section.
@@ 4 Fast Tensor Attention Gradient Computation
[…]
@@
13:  $U_{4}\leftarrow\operatorname{diag}(\widetilde{\mathsf{R}}(x))U_{1}$ $\triangleright$ $O(n^{1+o(1)})$ time
14:  $V_{4},~W_{4}\leftarrow V_{1},~W_{1}$ $\triangleright$ $O(n^{1+o(1)})$ time
@@
16:  $U_{5},~V_{5},~W_{5}\leftarrow[U_{3},-U_{4}],~[V_{3},V_{4}],~[W_{3},W_{4}]$ $\triangleright$ $O(n^{1+o(1)})$ time
@@
19:  $\widetilde{g}\leftarrow(A_{1}^{\top}U_{5})\odot(A_{2}^{\top}V_{5})\odot(A_{3}^{\top}W_{5})$ $\triangleright$ $\odot$ in Definition C.3. $O(d^{3}n^{o(1)})$ time
20:  return $\widetilde{g} \triangleright$ As $d=O(\log n)$, the total complexity is $O(n^{1+o(1)})$ time
@@ 4.1 Main Results for Fast Gradient Computation
Theorem 4.2 (Main result for fast gradient computation, informal version of Theorem F.8).
Assuming the entries of $A_{1},A_{2},$ $A_{3},$ $A_{4},A_{5},E\in\mathbb{R}^{n\times d}$ and $X_{1},X_{2},X_{3},Y_{1},Y_{2}\in\mathbb{R}^{d\times d}$ are represented using $O(\log n)$ bits. Then, there exist an algorithm (Algorithm 1) that runs in $n^{1+o(1)}$ time to solve $\mathsf{ATAttLGC}(n,d=O(\log n),B=o(\sqrt[3]{\log n}),\epsilon=1/\operatorname{poly}(n))$ (see Definition 2.9), i.e., our algorithm computes a gradient matrix $\widetilde{g}\in\mathbb{R}^{d\times d^{2}}$ satisfying
$\|\frac{\mathrm{d}\mathsf{Loss}(X)}{\mathrm{d}X}-\widetilde{g}\|_{\infty}\leq 1/\operatorname{poly}(n).$
@@
The assumption in Theorem 4.2 is practical. In practice, especially in recent long context tasks, the $n$ is large, e.g., $n=2\times 10^{6}$ for Google’s Gemini 1.5 Pro  [28], while the model training uses a half-precision floating-point format, e.g., the bit number is $16$. Furthermore, our assumption is “tight”, where if we slightly weaken the assumption, there is no algorithm that can solve the tensor attention gradient computation in truly sub-cubic complexity (Theorem 5.3).
@@ 4.2 Tensor Operation Analysis Techniques
We first introduce a distributed rule, where the proof is in Appendix C.2.
Fact 4.4.
@@
$~\underbrace{(U_{1}\ominus U_{2})}_{n_{1}\times dk}\cdot(\underbrace{(V_{1}\ominus V_{2})}_{n_{2}\times dk}\oslash\underbrace{(W_{1}\ominus W_{2})}_{n_{3}\times dk})^{\top}$
$=$ $~(\underbrace{U_{1}}_{n_{1}\times d}(\underbrace{V_{1}}_{n_{2}\times d}\oslash\underbrace{W_{1}}_{n_{3}\times d})^{\top})\circ(\underbrace{U_{2}}_{n_{1}\times k}(\underbrace{V_{2}}_{n_{2}\times k}\oslash\underbrace{W_{2}}_{n_{3}\times k})^{\top})$
Fact 4.4 tells us that the multiple tensor operation can be distributed to a different format. If we have some low-rank matrix/tensor, we can distribute them into each component so that each component can be accelerated via the low-rank property. Intuitively, this allows us to borrow some low-rank benefits from other terms to fix the bottleneck terms.
Then, we provide an important tool whose proof is in Appendix C.2.
@@
Fact 4.6 (Tensor-trick).
@@ 5.1 $\mathsf{SETH}$ and Tensor Attention Forward Hardness
Hypothesis 5.1 (Strong Exponential Time Hypothesis ($\mathsf{SETH}$), [35]).
@@
Assuming $\mathsf{SETH}$, for any constant $\delta>0$, no algorithm can solve $\mathsf{ATAttC}(n,d=\Theta(\log n),B=\Theta(\sqrt[3]{(1+\gamma)\log n}),\epsilon=n^{\gamma-O(1)})$ (Definition 2.7) in $O(n^{3-\delta})$ time, even if the inputs meet the following conditions for any $\gamma\geq 0$: (1) $V\in\{0,1\}^{n^{2}\times d}$, (2) There exists $B_{a}\leq O((1+\gamma)\log^{2}n)=O(d(\sqrt[3]{(1+\gamma)\log n})^{3})$ where all entries of $Q(K_{1}\oslash K_{2})^{\top}$ are within the range $1,B_{a}]$ and more than half entries in each row of $Q(K_{1}\oslash K_{2})^{\top}$ are equal to $B_{a}$.
This result shows that assuming $\mathsf{SETH}$, if we just slightly weaken the assumption from $B=O(\sqrt[3]{\log n})$ to $B=\Theta(\sqrt[3]{(1+\gamma)\log n})$ with $\gamma=\omega(1)$, then the tensor attention forward computation is hard, i.e., no algorithm can solve it in truly sub-cubic time.
@@ 5.2 Main Result for Hardness
Theorem 5.3 (Main result for hardness, informal version of Theorem G.3).
@@ 6 Discussion and Conclusion
In this work, we proved that the backward gradient of tensor attention training can be computed in almost linear $n^{1+o(1)}$ time, the same complexity as its forward computation, under a bounded entries assumption. We provided a closed-form solution for the gradient and proposed a fast computation method utilizing polynomial approximation and tensor algebraic techniques. Furthermore, we proved the necessity and tightness of our assumption through hardness analysis, showing that slightly weakening it renders the tensor attention gradient problem unsolvable in truly subcubic time.
Our theoretical results establish the feasibility of efficient higher-order transformer training and may facilitate practical applications of tensor attention architectures. Due to space limits, we provide our further discussion and extension in Appendix A. Future work can perform empirical evaluations of the method in practical large language models, and explore how these findings can be implemented in real-world scenarios to enable the development of powerful higher-order models.
@@
In Section A, we provide a further discussion and extension of this work. In Section B, we provide related works. In Section C, we provide general definitions and several basic facts. In Section D, we show how we calculate the gradient of the loss function. In Section E, we show the time complexity of our algorithm. In Section F, we show that our algorithm can be computed in polynomial time. In Section G, we show the hardness of our algorithm. In Section LABEL:sec:lim, we discuss the limitation of this work. In Section LABEL:sec:impact, we provide an elaborate discussion about potential societal impacts.
@@ Appendix A Further Discussion and Extension
Technical novelty over previous works. We generalize beyond the results of [7], which only provide methods for tensor attention forward. Our paper presents a detailed analysis for tensor attention backward, providing both upper bound and lower bound. Though we build on some results from [7] and [8], generalizing to tensor attention backward posed many technical challenges. These challenges are unique to our setting and not presented in previous settings like matrix attention [6, 8] or tensor attention forward [7]. To be more specific, we prove many key properties for the tensor operation needed for backward though not needed for forward, including 4.4 (distribution rule for tensor and matrix product), C.11 (tensor computation reduction to matrix product), C.12 (distribution rule for tensor computation), Claim C.20 (tensor product to matrix product). Lemma 4.5 supports the proof of Fact 4.4 and helps bypass the $O(n^{3}d^{2})$ time complexity bottleneck in the fast computation of $U_{2}$. Fact 4.4, crucial in proving Lemma F.5, shows the distributive nature of tensor operations. Using Facts C.11, C.12, and Claim C.20, we leverage the structure of low-rank matrices $U_{5},V_{5},W_{5}$ to prove Theorem 4.2.
@@
Appendix B Related Work
Fast attention computation.
In recent years, significant advances have been made in the development of efficient attention computation. One research direction involves employing low-rank approximations, polynomial kernel, or random features for the attention matrix [19, 100, 6, 37, 77, 30], which scales the computational complexity sub-quadratically with sequence length. Another method explores patterns of sparse attention that lessen the computational load [32]. Additionally, using linear attention as an alternative to softmax attention has emerged as a substantial area of study [40, 75, 98, 3, 96]. These innovations have enhanced the capability of transformer-based models to handle longer sequences, thereby broadening their potential applications across various fields [18, 78, 67, 23, 57, 12, 36]. On the other hand, FlashAttention [20, 21] is one of the most popular practical methods to accelerate attention computation, while it achieves a considerable improvement in running time with a constant complexity ratio.
Tensor computation for high-order representation.
Tensors excel over matrices in capturing higher-order relationships within data [99]. Calculating low-rank factorizations or approximations of tensors is essential in a wide range of computer science applications, such as natural language processing [45, 14], computer vision [54, 17], computer graphics [88, 86], security [1, 42], and data mining [38, 72, 64]. Moreover, tensors are crucial in numerous machine learning applications [68, 101, 93] and other diverse fields [73, 94, 71].
@@ Appendix C Tensor Operation Background
In Section C.1, we define the notation of computational time and the tensor operation. In Section C.2, we provide some helpful facts of tensor operation. In Section C.3, we provide some helpful facts of vectorization operation. In Section C.4, we provide some helpful facts about the tensor product. It is worth noting that proofs for some of the facts discussed in this section are also available in [43].
@@ C.1 General definitions and tensor operation
Definition C.2 (Third mode tensor product $(\cdot,\cdot,\cdot)$).
@@ C.2 Facts for tensor operation
In Fact C.5, due to definition $V_{1}$ and $V_{2}$ need to have the same number of rows. $W_{1}$ and $W_{2}$ also need to have the same number of rows. $V_{1}$ and $W_{1}$ need to have same number of columns, and $V_{2}$ and $W_{2}$ need to have same number of columns.
Fact C.7 (Swap rule for tensor product and matrix product).
@@
Fact C.8 (Restatement of Fact 4.4).
@@
where the first step follows from Definition C.2, the second step follows from the property of identity tensor $(\mathsf{I}_{d})_{i,j,k}$, which equals 11 only when $i=j=k$ and 0 elsewhere, and the last step follows from Definition C.3. ∎
@@
where the first step follows from definition of $C\in\mathbb{R}^{d\times d}$, the second step follows from the matrix can written as the summation of $n_{1}n_{2}$ rank-11 matrices, the third step follows from changing the index, the forth step follows from $\underbrace{A_{i+(j-1)n_{1},*}}_{d_{1}\times 1}=\underbrace{a_{1,i}}_{d_{1}\times 1}\circ\underbrace{a_{2,j}}_{d_{1}\times 1}$ by Definition 2.2.
@@ C.3 Facts for vectorization operation
[…]
@@
Fact C.16 (Tensor-trick, Restatement of Fact 4.6).
@@ C.4 Facts for tensor product
Let $\mathsf{I}_{d}\in\mathbb{R}^{d\times d\times d}$ and $A_{1},A_{2},A_{3}\in\mathbb{R}^{n\times d}$. We have $\mathrm{mat}(\mathsf{I}_{d}(A_{1},A_{2},A_{3}))=A_{1}\mathrm{mat}(\mathsf{I}_{d})(A_{2}\otimes A_{3})^{\top}=A_{1}(A_{2}\oslash A_{3})^{\top}\in\mathbb{R}^{n\times n^{2}}$ .